# Examining the Expanding Role of Synthetic Data Throughout the AI Development Pipeline

### v1

@@ 4.3. Validating Synthetic Data
Approaches to data validation. Practitioners described the validation of synthetic data as “subtle work” (P18) that was subjective and required them to rely on their intuition. Manual inspection—also referred to as “spot-checking” or “eyeballing” by participants—emerged as the most common validation method. P3 described a periodic monitoring routine in which they inspected examples every few weeks to “make sure the data is as it should be, and that things are going smoothly.” During these inspections, they often discovered cases where the model “was not following the prompt correctly and things don’t make sense.” Similarly, P5, who focused on measuring harms, shared their process of manually reviewing datasets to check for alignment with instructions, explaining “I would generate a dataset and then scroll through it, asking, ‘Does this make sense? Is it following the instructions we’ve given it?’” Practitioners acknowledged that this approach made it challenging to maintain consistency, especially when the data was inspected by different team members.
@@
Alternative methods beyond manual inspection were more rare. P4, whose work involved simulating user data, described how they “don’t yet have a methodology for evaluation” of synthetic data. Instead, they focused on the usefulness of the data in downstream applications. A few, including P15 and P18, conducted small-scale studies with human experts (e.g., doctors or trust and safety professionals) to validate synthetic data for specialized tasks. These participants also performed ablation experiments by generating small synthetic datasets and evaluating their impact on the primary model’s performance when used for training. P7 described how their team used the internal confidence score of the auxiliary model in validating its own output, noting the circular logic. They set thresholds to classify outputs, treating synthetic data with confidence scores above 90% as high quality and those around 30% as low quality.
@@
Challenges in data validation. Participants described validation of synthetic data as a consistent bottleneck. Ironically, the qualities that made synthetic data attractive–such as its scale, diversity, and the ability to simulate data for rare scenarios—also made it exceptionally challenging to validate since, as described above, validation was typically manual.
@@
Certain types of data were especially challenging to inspect manually. P8, who worked in medical imaging, described the challenges of confirming whether synthetic images accurately reflected specific medical conditions, particularly those with subtle visual markers. They collaborated with a resident doctor to assess the realisticness of their data. Practitioners working in high-stakes domains, like harmful content detection or red-teaming, encountered similar challenges.
@@
Another significant challenge was defining what qualifies as ‘good’ synthetic data, especially in contexts where no real data exists to guide conceptualization. As described above, practitioners sought data that appeared diverse and natural-looking. Yet, as many participants noted, it was difficult to articulate what these criteria mean. For example, P5, who focused on establishing safety evaluations for generative AI products, described how their workflow lacked user data to help gauge the diversity and distribution of model outputs. Practitioners often defaulted to defining diversity as the generation of non-redundant examples, such as P16 who described how “the main way [they] defined diversity was how many unique examples could be generated. If [they] call the generator model a thousand times, how many redundant examples were generated?” However, simply avoiding redundancy did not guarantee that the data covers a meaningful range of variation. P3, who focused on detecting stereotypes in primary model outputs, found it challenging to define the types of stereotypes they aimed to capture and operationalize these criteria within prompts. And P17 struggled to validate whether the synthetic data accurately reflected underrepresented groups, noting the lack of clear indicators to confirm that the auxiliary model was following instructions or capturing the intended range of diversity.
@@
Lastly, many practitioners highlighted that their companies prioritized scaling data generation to keep pace with model development demands, relegating validation to a lower priority. P3 expressed this tension: “Honestly, people just don’t want to do [the validation] right now because of the urgency associated with everything.” While this participant recognized the risks of insufficient validation, they felt pressured to move forward to meet internal deadlines, hoping to address data quality issues later if time and resources allowed.
@@ 4.4. General Limitations and Ethical Considerations of Synthetic Data
‘Chaining’ auxiliary models. Many practitioners described using the same or similar auxiliary models to produce both training and evaluation data, or to generate evaluation data and score the primary model’s outputs. This practice of ‘chaining’ auxiliary models raised concerns about the risk of amplifying biases at different stages of the AI development pipeline. P10 gave a hypothetical example where the auxiliary model was the same as the model being evaluated: “GPT-4 as a scorer and asking it to evaluate a GPT-4 based prediction and comparing that to a LLaMA prediction. There’s a high chance GPT-4 based prediction will be scored higher because your scoring module is also GPT-4.” E24 reinforced how the cumulative effect of seemingly “minor problems or morally neutral outcomes” could become significant when repeated on a massive scale: “You’ve done something that maybe is not that bad, but you have done it 10,000 times instead of 100 times, and you have also done it in a way where you are not really sure what the gaps are and what the dangers are.” They argued that chaining models creates feedback loops that mask issues that would emerge through more diverse evaluation methods.
@@
Removed avenues for exercising agency. RAI experts brought up that the use of synthetic data undermines stakeholders’ ability to exert meaningful agency over AI systems. They emphasized that synthetic data creates distance between individuals and the data that was originally derived from their activities, decisions, or behaviors. This layer of abstraction erodes opportunities to contest the ways that datasets are used. As E24 noted, “the link between a person and their data and a training dataset” is arguably one of the most effective opportunities for participation or redress.
Several RAI experts remarked on the extractive nature of generating synthetic data using auxiliary models that may have been developed through the expropriation of the work of data subjects. E27 linked these concerns to ongoing debates in the digital humanities, where artists, writers, and creators have called for ethical engagement with data derived from their work, critiquing practices that commodify their intellectual and creative output (cf. Jiang et al., 2023). E28 suggested that the purported privacy-preserving benefits of synthetic data often come at the expense of others. They explained that for auxiliary models to generate data about a specific group, they require relevant information in the training data: “so how good this [data is] is a matter of how well it is violating someone else’s privacy, right?”
Stereotyping and cultural hegemony. Many participants highlighted risks of using identity-related prompting (i.e., including a series of adjectives to describe a simulated user’s identity) when generating data. E25 noted that this practice could cause an auxiliary model to either produce stereotypical patterns of behavior or generate overly generic responses due to its alignment process. P6 and E26 both noted that synthetic data intended to reflect identity often lacked nuance and depth, in part because auxiliary models rely on data about a group rather than data generated by members of the group (cf. Wang et al., 2024). As P6 put it: “Most of that data is in a tone or a stance, or a perspective of being about these different groups of people, and not by these groups of people, and so there’s a lot of cases where we want to generate data as a certain demographic, or as a certain marginalized group, and it almost feels like a caricature of that group.
@@
Organizational pressure to scale. While practitioners acknowledged that synthetic data has its limitations, many expressed a sense of resignation toward addressing them. Practitioners described how they were navigating a fast-paced and competitive AI landscape where organizational priorities often emphasized scaling data generation to meet the demands of model development and evaluation. While practitioners recognized that synthetic data could introduce biases or fail to capture the complexity of real-world data, they frequently proceeded despite these risks, with the intention of addressing potential issues later. P18 recognized this recurring tension: “Frankly, this is a very busy space. I would love to take things slower, but it’s just– I don’t know. We’ve had multiple failed attempts to do better documentation of our synthetic data, have better practices, which did not go very well with how fast we worked honestly.
@@ 5. Discussion
While our findings highlight significant concerns about the role of auxiliary models in development practices, we resist framing synthetic data as inherently problematic. An AI development pipeline that excludes synthetic data may involve other challenges, for instance, in meeting the ever-increasing demands of evaluation for robust model development across a range of harm areas and specialized domains, as required by policymakers or industry standards. We contend that the impacts of synthetic data are multifaceted and plural, with both potential benefits and risks, and invite critical, ongoing investigation by the research community.
@@ 5.1. Interrogating Impacts to the AI Supply Chain
Our findings capture the perspectives of practitioners embedded within systems of institutional power. This ‘studying up’ approach (Nader, 1972; Barabas et al., 2020) may fall short in surfacing how synthetic data use reconfigures other aspects of the AI supply chain (Widder and Wong, 2023). We identify opportunities to investigate impacts on model producers, data workers, and model subjects.
Building AI from scratch requires vast amounts of high-quality data that raise the barrier for entry for new companies seeking to develop applications (Widder et al., 2023a). One might argue synthetic data could contribute to the democratization of model development since smaller companies no longer need to have access to large-scale, costly, hard-to-obtain datasets to train their models (Lee, 2024). Synthetic data could counteract the concentration of power within the hands of a few incumbents by allowing startups or other entities to compete against large tech companies.
Nevertheless, there is a paradox in this democratization narrative. The auxiliary models typically used to generate synthetic data are almost exclusively developed by the same tech giants that already dominate the AI industry (Widder et al., 2023a). Smaller companies or researchers who wish to train a model themselves are likely to rely on these auxiliary models offered as-a-service to produce synthetic data (Luitse and Denkena, 2021). This results in a consolidation of power as big tech companies dictate the tools and frameworks that smaller companies must rely on for generating synthetic data (Azoulay et al., 2024).
@@
On the other hand, auxiliary model use offers potential opportunities to mitigate some of the harms associated with traditional data annotation, particularly in cases where workers are exposed to toxic or harmful content. Participants mentioned how automating the evaluation of certain types of harms (e.g., those involving violent, graphic content) and relying on annotators for subtle or complex forms of harm could reduce the need for human annotators to engage directly with harmful material that has been linked to psychological distress and trauma (Steiger et al., 2021).
@@
Finally, the use of synthetic data for training and evaluating models has significant implications for model subjects. Studies have shown that distribution shifts caused by synthetic data during training can lead to reductions in model performance and fairness and to increased representational harms for minoritized groups (Wyllie et al., 2024). If auxiliary models are also used in the evaluation, they might fail to detect critical issues, particularly those that disproportionately affect underrepresented groups or rare phenomena. This risk of degraded model performance during training is likely unnoticed if evaluation also highlights the urgency of advancing research in this area. What mechanisms could then be developed to scrutinize and mitigate the ways in which general-purpose models propagate biases into synthetic datasets?
5.2. Towards Considerations for Responsible Use of Synthetic Data
@@
Our interviews reflect a particular moment in time when practices and norms around the use of synthetic data in both industry and research contexts are rapidly evolving. The viewpoints expressed by different participants were often in tension with each other or with the growing synthetic data literature (Qian et al., 2024). Some participants argued that synthetic data is well-suited for use in training due to tolerance for noise but too risky to use in evaluation, while others expressed the opposite viewpoint. Some praised the use of synthetic data because it gave them “full control,” while others recounted difficulties controlling data generation in practice. Some proposed that auxiliary models can be used to simulate data from members of underrepresented groups, a viewpoint some literature supports (Grossmann et al., 2023), while others argued that data generated in this way would be merely a “caricature” of the group, in line with other scholarship (Wang et al., 2024). These disagreements reflect the contested nature of our collective understanding of synthetic data’s benefits, limitations, and appropriate use. Because of this, rather than offering prescriptive recommendations, we focus on foregrounding key considerations that researchers and practitioners must navigate as they engage with synthetic data.
@@
Practitioners should carefully consider what constitutes an appropriate application of synthetic data. Deciding whether to use synthetic data and in what capacity involves assessing the intended purpose or use case and the domain of application and balancing trade-offs between resource availability and the ‘right’ methodology for data generation and validation. For example, in a ‘high-risk’ domain (like training a model for a medical application), it is arguably more crucial to clearly operationalize the desiderata (diversity, for instance) and to systematically validate that the data meets those characteristics. In ‘low-risk’ [^2] What counts as high-risk or low-risk is a value-laden assumption that must itself be interrogated. domains, there is perhaps more tolerance for noise or fuzzy data. Future research could investigate which dimensions of a use case or domain are relevant in determining the appropriate scope of use for synthetic data.
Synthetic data practices are heavily entangled with selection of auxiliary models. The harms of synthetic data are a result of the assumptions and design decisions during the creation of synthetic data and the limitations of the auxiliary models. Each model has a different risk profile (Uuk et al., 2024), which makes auxiliary model selection a consequential decision. Despite this, participants in our study often defaulted to selecting what they perceived as the state-of-the-art model.
@@
This tendency reflects a broader pattern in AI practices, where newer models are assumed to outperform older ones (Wong, 2024), even when their suitability for particular contexts remains untested. Practitioners could benefit from a more deliberate approach to model selection where they experiment with whether a particular model aligns with their specific needs.
@@
In addition to careful model selection, there is a need to systematize synthetic data validation. Participants in this study relied on ‘eye-balling’ or ‘spot-checking’ as their primary validation method. While manual review offers a direct way to identify errors, participants frequently reported it as inconsistent, insufficient, or deprioritized over demands of data generation. First, practitioners need to invest more time and resources in synthetic data validation. Second, practitioners should develop a methodology that ideally combines multiple forms of data validation (e.g., manual verification and calibration experiments). There would be several considerations to keep in mind. For example, instead of randomly sampling data, one could prioritize defining (including with the absence of real-world data) and creating underrepresented or edge-case examples. Similarly, to guide validation efforts, practitioners could develop a rubric that articulates the criteria against which they compare the synthetic data. Yet, how can such qualities be effectively measured, and how might reliance on proxy metrics introduce additional risks? We call for research and experimentation to develop effective validation practices for synthetic data.
@@
Lastly, AI workflows that incorporate auxiliary models must prioritize meaningful community engagement. Synthetic data inherently reduces opportunities for direct participation, as it is neither created nor annotated by individuals or communities. Often, an underlying motivation for using synthetic data is to “take the people out” (Susser et al., 2024a), whether to address privacy concerns, meet compliance requirements, or minimize exposure to harmful content. While these intentions may serve some goals, they also exclude data subjects and data workers from the pipeline and restrict opportunities to question or shape datasets. To increase participation, participants in our study proposed a consultation model where experts or community members provide input on use cases, evaluation outcomes, or critical decisions (Delgado et al., 2023). However, when is it appropriate to consult only experts, and when should affected communities play a central role? The answer likely depends on the specific use case and the degree of transformation involved. For example, small or targeted modifications to existing datasets might raise different participatory needs compared to entirely synthetic datasets. The timing of participation is equally significant. Should stakeholders contribute during the generation process, shaping decisions about what data is created and how? Or is it more effective to involve them during validation? These decisions are not merely procedural, but shape the accountability of synthetic data practices to the communities they may impact.
5.3. Potential adverse impacts.
Our research has several potential adverse impacts. First, our findings focus exclusively on practitioners’ and RAI experts’ perspectives. Our findings may inadvertently privilege their viewpoints and concerns over those of other crucial stakeholders. This could lead to recommendations that undervalue the needs of affected communities, policymakers, and civil society organizations. Second, our research sheds light on motivations, practices, and challenges for the use of synthetic data in the AI development pipeline. We also expose limitations and ethical considerations. Emphasizing these drawbacks may be interpreted as evidence that the use of synthetic data should be more restricted. However, we caution against this interpretation of our results. In setting policy guidance for synthetic data, policymakers should weigh the benefits as well. Perhaps most notably, synthetic data enables the evaluation of models and systems at a scale that would not be possible through other means; without it, many evaluations simply may not happen. To address this concern, we have aimed throughout this paper to provide a balanced perspective, highlighting both the benefits and drawbacks of synthetic data. We additionally emphasize that norms and practices around the use of synthetic data are rapidly evolving. While we hope that our research inspires future work developing best practices for the responsible use of synthetic data, these practices will need to be constantly evaluated and updated to ensure that they remain practical and relevant.

### v2

@@ 4.3. Validating Synthetic Data
Practitioners described the validation of synthetic data as “subtle work” (P18) that was subjective and required them to rely on their intuition. Manual inspection—also referred to as “spot-checking” or “eyeballing” by participants—emerged as the most common validation method. P3 described a periodic monitoring routine in which they inspected examples every few weeks to “make sure the data is as it should be, and that things are going smoothly.” During these inspections, they often discovered cases where the model “was not following the prompt correctly and things don’t make sense.” Similarly, P5, who focused on measuring harms, shared their process of manually reviewing datasets to check for alignment with instructions, explaining “I would generate a dataset and then scroll through it, asking, ‘Does this make sense? Is it following the instructions we’ve given it?’” Practitioners acknowledged that this approach made it challenging to maintain consistency, especially when the data was inspected by different team members.
@@
Alternative methods beyond manual inspection were rarer. P4, whose work involved simulating user data, described how they “don’t yet have a methodology for evaluation” of synthetic data. Instead, they focused on the usefulness of the data in downstream applications. A few, including P15 and P18, conducted small-scale studies with human experts (e.g., doctors or trust and safety professionals) to validate synthetic data for specialized tasks. These participants also performed ablation experiments by generating small synthetic datasets and evaluating their impact on the primary model’s performance when used for training. P7 described how their team used the internal confidence score of the auxiliary model in validating its own output, noting the circular logic. They set thresholds to classify outputs, treating synthetic data with confidence scores above 90% as high quality and those around 30% as low quality.
4.3.2. Challenges in Data Validation.
@@
Participants described validation of synthetic data as a consistent bottleneck. Ironically, the qualities that made synthetic data attractive–such as its scale, diversity, and the ability to simulate data for rare scenarios–also made it exceptionally challenging to validate since, as described above, validation was typically manual.
@@
Certain types of data were especially challenging to inspect manually. P8, who worked in medical imaging, described the challenges of confirming whether synthetic images accurately reflected specific medical conditions, particularly those with subtle visual markers. They collaborated with a resident doctor to assess the realism of their data. Practitioners working in high-stakes domains, like harmful content detection or red-teaming, encountered similar challenges.
@@
Another significant challenge was defining what qualifies as ‘good’ synthetic data, especially in contexts where no real data exists to guide conceptualization. As described above, practitioners sought data that appeared diverse and natural-looking. Yet, as participants noted, it was difficult to articulate what these criteria mean. For example, P5, who focused on establishing safety evaluations for generative AI products, described how their workflow lacked user data to help gauge the diversity and distribution of model outputs. Practitioners often defaulted to defining diversity as the generation of non-redundant examples, such as P16, who described how “the main way [they] defined diversity was how many unique examples could be generated. If [they] call the generator model a thousand times, how many redundant examples were generated?” However, simply avoiding redundancy did not guarantee that the data covered a meaningful range of variation. P3, who focused on detecting stereotypes in primary model outputs, found it challenging to define the types of stereotypes they aimed to capture and operationalize these criteria within prompts. And P17 struggled to validate whether the synthetic data accurately reflected underrepresented groups, noting the lack of clear indicators to confirm that the auxiliary model was following instructions or capturing the intended range of diversity.
@@
Lastly, practitioners highlighted that their companies prioritized scaling data generation to keep pace with model development demands, relegating validation to a lower priority. P3 expressed this tension: “Honestly, people just don’t want to do [the validation] right now because of the urgency associated with everything.” While this participant recognized the risks of insufficient validation, they felt pressured to move forward to meet internal deadlines, hoping to address data quality issues later if time and resources allowed.
@@ 4.4. General Limitations and Ethical Considerations of Synthetic Data
4.4.1. ‘Chaining’ Auxiliary Models.
Practitioners described using the same or similar auxiliary models to produce both training and evaluation data, or to generate evaluation data and score the primary model’s outputs. This practice of ‘chaining’ auxiliary models raised concerns about the risk of amplifying biases at different stages of the AI development pipeline. P10 gave a hypothetical example where the auxiliary model was the same as the model being evaluated: “GPT-4 as a scorer and asking it to evaluate a GPT-4 based prediction and comparing that to a LLaMA prediction. There’s a high chance GPT-4-based prediction will be scored higher because your scoring module is also GPT-4.” E24 reinforced how the cumulative effect of seemingly “minor problems or morally neutral outcomes” could become significant when repeated on a massive scale: “You’ve done something that maybe is not that bad, but you have done it 10,000 times instead of 100 times, and you have also done it in a way where you are not really sure what the gaps are and what the dangers are.” They argued that chaining models creates feedback loops that mask issues that would emerge through more diverse evaluation methods.
@@
4.4.2. Removed Avenues for Exercising Agency.
RAI experts brought up that the use of synthetic data undermines stakeholders’ ability to exert meaningful agency over AI systems. They emphasized that synthetic data creates distance between individuals and the data that was originally derived from their activities, decisions, or behaviors. This layer of abstraction erodes opportunities to contest the ways that datasets are used. As E24 noted, “the link between a person and their data and a training dataset” is arguably one of the most effective opportunities for participation or redress.
Several RAI experts remarked on the extractive nature of generating synthetic data using auxiliary models that may have been developed through the expropriation of the work of data subjects. E27 linked these concerns to ongoing debates in the digital humanities, where artists, writers, and creators have called for ethical engagement with data derived from their work, critiquing practices that commodify their intellectual and creative output (cf. Jiang et al., 2023). E28 suggested that the purported privacy-preserving benefits of synthetic data often come at the expense of others, explaining that for auxiliary models to generate data about a specific group, they require relevant information in the training data: “so how good this [data is] is a matter of how well it is violating someone else’s privacy, right?”
4.4.3. Stereotyping and Cultural Hegemony.
Participants highlighted risks of using identity-related prompting (i.e., including a series of adjectives to describe a simulated user’s identity) when generating data. E25 noted that this practice could cause an auxiliary model to either produce stereotypical patterns of behavior or generate overly generic responses due to its alignment process. P6 and E26 both noted that synthetic data intended to reflect identity often lacked nuance and depth, in part because auxiliary models rely on data about a group rather than data generated by members of the group (cf. Wang et al., 2024). As P6 put it: “Most of that data is in a tone or a stance, or a perspective of being about these different groups of people, and not by these groups of people, and so there’s a lot of cases where we want to generate data as a certain demographic, or as a certain marginalized group, and it almost feels like a caricature of that group.
@@
4.4.4. Organizational Pressure to Scale.
While practitioners acknowledged that synthetic data has its limitations, many expressed a sense of resignation toward addressing them. Practitioners described how they were navigating a fast-paced and competitive AI landscape where organizational priorities often emphasized scaling data generation to meet the demands of model development and evaluation. While practitioners recognized that synthetic data could introduce biases or fail to capture the complexity of real-world data, they frequently proceeded despite these risks, with the intention of addressing potential issues later. P18 recognized this recurring tension: “Frankly, this is a very busy space. I would love to take things slower, but it’s just– I don’t know. We’ve had multiple failed attempts to do better documentation of our synthetic data, have better practices, which did not go very well with how fast we worked honestly.
@@ 5. Discussion
While our findings highlight significant concerns about the role of auxiliary models in development practices, we resist framing synthetic data as inherently problematic. An AI development pipeline that excludes synthetic data may involve other challenges, for instance, in meeting the ever-increasing demands of evaluation for robust model development across a range of harm areas and specialized domains, as required by policymakers or industry standards. We contend that the impacts of synthetic data are multifaceted and plural, with both potential benefits and risks, and invite critical, ongoing investigation by the FAccT community.
@@ 5.1. Interrogating Impacts to the AI Supply Chain
Our findings capture the perspectives of practitioners embedded within systems of institutional power. This ‘studying up’ approach (Nader, 1972; Barabas et al., 2020) may fall short in surfacing how synthetic data use reconfigures other aspects of the AI supply chain (Widder and Wong, 2023). Below, we outline opportunities to investigate impacts on model deployers, data workers, and model subjects.
Building AI from scratch requires vast amounts of high-quality data that raises the barrier for entry for new companies seeking to develop applications (Widder et al., 2023a). One might argue that synthetic data could contribute to the democratization of model development since smaller companies no longer need to have access to large-scale, costly, hard-to-obtain datasets to train their models (Lee, 2024). In theory, synthetic data could counteract the concentration of power within the hands of a few incumbents by allowing startups or other entities to compete against large tech companies.
Nevertheless, there is a paradox in this democratization narrative. The auxiliary models typically used to generate synthetic data are almost exclusively developed by the same tech giants that already dominate the AI industry (Widder et al., 2023a). Smaller companies or researchers who wish to train a model themselves are likely to rely on these auxiliary models offered as-a-service to produce synthetic data (Luitse and Denkena, 2021). This results in a consolidation of power as big tech companies monopolize the tools and frameworks (Burkhardt and Rieder, 2024) that smaller companies must rely on for generating synthetic data (Azoulay et al., 2024). As synthetic data pipelines become entangled with proprietary model APIs, usage restrictions, and tiered access to compute, the ability to experiment or contest how synthetic data is generated is increasingly constrained.
@@
On the other hand, auxiliary model use offers potential opportunities to mitigate some of the harms associated with traditional data annotation, particularly in cases where workers are exposed to toxic or harmful content. Participants mentioned how automating the evaluation of certain types of harms (e.g., those involving violent, graphic content) and relying on annotators for subtle or complex forms of harm, could reduce the need for human annotators to engage directly with harmful material that has been linked to psychological distress and trauma (Steiger et al., 2021).
@@
Finally, the use of synthetic data for training and evaluating models has significant implications for model subjects. Studies have shown that distribution shifts caused by synthetic data during training can lead to reductions in model performance and fairness, and to increased representational harms for minoritized groups (Wyllie et al., 2024). If auxiliary models are also used in the evaluation, they might fail to detect critical issues, particularly those that disproportionately affect underrepresented groups or rare phenomena. This risk of degraded model performance during training, which is likely unnoticed in evaluation, also highlights the urgency of advancing research in this area. What mechanisms could then be developed to scrutinize and mitigate the ways in which general-purpose models propagate biases into synthetic datasets?
5.2. Toward Considerations for Responsible Use of Synthetic Data
@@
Our interviews reflect a particular moment in time when practices and norms around the use of synthetic data in both industry and research contexts are rapidly evolving. The viewpoints expressed by different participants were often in tension with each other or with the growing synthetic data literature (Qian et al., 2024). Some participants argued that synthetic data is well-suited for use in training due to its tolerance for noise but is too risky to use in evaluation, while others expressed the opposite viewpoint. Some praised the use of synthetic data because it gave them “full control,” while others recounted difficulties controlling data generation in practice. Some proposed that auxiliary models can be used to simulate data from members of underrepresented groups, a viewpoint some literature supports (Grossmann et al., 2023), while others argued that data generated in this way would be merely a “caricature” of the group, in line with other scholarship (Wang et al., 2024). These disagreements reflect the contested nature of our collective understanding of synthetic data’s benefits, limitations, and appropriate use. Because of this, rather than offering prescriptive recommendations, we focus on foregrounding key considerations that researchers and practitioners must navigate as they engage with synthetic data.
@@
Practitioners should carefully consider what constitutes an appropriate application of synthetic data. Deciding whether to use synthetic data, and in what capacity, involves assessing the intended purpose or use case and the domain of application, and balancing trade-offs between resource availability and the ‘right’ methodology for data generation and validation. For example, in a ‘high-risk’ domain (like training a model for a medical application), it is arguably more crucial to clearly operationalize the desiderata (diversity, for instance) and to systematically validate that the data meets those characteristics. In ‘low-risk’ domains, there is perhaps more tolerance for noise or fuzzy data. Future research could investigate which dimensions of a use case or domain are relevant in determining the appropriate scope of use for synthetic data. Of course, what counts as high-risk or low-risk is itself a value-laden assumption that must be interrogated.
@@
Practitioners should carefully consider auxiliary model capabilities and limitations when selecting which model to use. Synthetic data practices are heavily entangled with the selection of auxiliary models. The harms of synthetic data are a result of the assumptions and design decisions made during the creation of synthetic data and the limitations of the auxiliary models. Each model has a different risk profile (Uuk et al., 2024), which makes auxiliary model selection a consequential decision. Despite this, participants in our study often defaulted to selecting what they perceived as the state-of-the-art model. This tendency reflects a broader pattern in AI practices, where newer models are assumed to outperform older ones (Wong, 2024), even when their suitability for particular contexts remains untested. Practitioners could benefit from a more deliberate approach to model selection, where they experiment with whether a particular model aligns with their specific needs.
@@
In addition to careful model selection, there is a need to make synthetic data validation more systematic. Participants in this study relied on ‘eye-balling’ or ‘spot-checking’ as their primary validation method. While manual review offers a direct way to identify errors, participants frequently reported it as inconsistent, insufficient, or deprioritized over the demands of data generation. First, practitioners need to invest more time and resources in synthetic data validation. Second, practitioners should develop a methodology that ideally combines multiple forms of data validation (e.g., manual verification and calibration experiments). There would be several considerations to keep in mind. For example, instead of randomly sampling data, one could prioritize defining (including with the absence of real-world data) and creating underrepresented or edge-case examples. Similarly, to guide validation efforts, practitioners could develop a rubric that articulates the criteria against which they compare the synthetic data. Yet, how can such qualities be effectively measured, and how might reliance on proxy metrics introduce additional risks? We call for research and experimentation to develop effective validation practices for synthetic data.
@@
Lastly, AI workflows that incorporate auxiliary models must prioritize meaningful community engagement. Synthetic data inherently reduces opportunities for direct participation, as it is neither created nor annotated by individuals or communities. Often, an underlying motivation for using synthetic data is to “take the people out” (Susser et al., 2024a), whether to address privacy concerns, meet compliance requirements, or minimize exposure to harmful content. While these intentions may serve some goals, they also exclude data subjects and data workers from the pipeline and restrict opportunities to question or shape datasets. To increase participation, participants in our study proposed a consultation model where experts or community members provide input on use cases, evaluation outcomes, or critical decisions (Delgado et al., 2023). However, when is it appropriate to consult only experts, and when should affected communities play a central role? The answer likely depends on the specific use case and the degree of transformation involved. For example, small or targeted modifications to existing datasets might raise different participatory needs compared to entirely synthetic datasets. The timing of participation is equally significant. Should stakeholders contribute during the generation process, shaping decisions about what data is created and how? Or is it more effective to involve them during validation? These decisions shape the accountability of synthetic data practices to the communities they may impact.
In highlighting these considerations, we recognize how meaningful change requires rethinking incentives for different stakeholders across the AI supply chain. First, regulation could establish reporting requirements for both auxiliary model producers and downstream developers. For model producers, this could involve disclosing evaluations of their models’ capacity to generate synthetic data in specific domains, along with guidance on the intended scope of use. Developers using synthetic data could be expected to document prompts, generation parameters, and any modifications made to the outputs. Furthermore, demonstrating that this kind of reporting leads to better outcomes (e.g., product usage or increased user trust) would establish clear incentives for these evaluations.